# Discrete Diffusion-Based Model-Level Explanation of Heterogeneous GNNs with Node Features

### v1

@@ Abstract
Introduction
@@
Figure 1: Overview of DiGNNExplainer. The example, based on the DBLP dataset, illustrates the process of generating explanations for a node classification task. Nodes represent different entity types: authors (red, $\bullet$), papers (blue, $\bullet$), conferences (orange, $\bullet$), and terms (green, $\bullet$). Node features, such as discrete bag-of-words representations of author keywords (e.g., [010100]), are encoded in red. The task is to predict the class of each author, i.e., their research area: database, data mining, artificial intelligence, or information retrieval. (1) & (2) The graph generator takes as input small real-world graphs with discrete node features and produces synthetic graphs with discrete node features. (3) The validity of these generated graphs is checked against the metagraph. (4) A trained GNN, which is to be explained, is applied to each of the nine author nodes of valid, synthesized graphs producing a prediction for each of the four classes. For each class, we identify the node and corresponding graph with the highest prediction probability, yielding one explanation graph per class. (5) To evaluate our approach, we compute various metrics, including the ground-truth faithfulness between our explanation graphs and the communities within the real dataset.
@@
Complex real-world data, such as citation networks, social networks, and molecular structures, often exhibit an underlying graph structure. Nodes in these graphs typically have different types and are associated with additional features, i.e., they are heterogeneous graphs. For instance, in a citation network, there might be nodes of type “Paper” and “Author” each of which has additional node features like keywords or affiliations. Whereas related work sometimes uses the term “node feature” only to refer to a one-hot encoding of the node type (Yuan et al. 2020; Wang and Shen 2023), we refer to actual attributes of nodes (e.g., age of persons, keywords of papers). A key machine learning task on such graphs is node classification, e.g., for fake news detection (Zhang et al. 2024a), corporate risk assessment (Zhang et al. 2024b), or predicting the mutagenicity of molecules (Luo et al. 2020). Towards this end, a number of powerful approaches for heterogeneous graphs have been proposed, e.g., Heterogeneous Graph Neural Networks (HGNNs)(Wang et al. 2019).
@@
However, the predictions made by HGNNs are opaque, providing no insight into the reasoning behind their decisions. In high-stakes applications—such as healthcare, finance, or criminal justice—this lack of transparency is a significant limitation, as it prevents users from understanding the rationale behind the model’s decisions. Towards this end, several post-hoc explanation methods have been proposed, such as GNNExplainer (Ying et al. 2019), SubgraphX (Yuan et al. 2021), PGExplainer (Luo et al. 2020), XGNN (Yuan et al. 2020), GNNInterpreter (Wang and Shen 2023), and D4Explainer (Chen et al. 2023). Despite their contributions, these methods face significant limitations: (1) the first three do not provide model-level explanations, (2) the second three lack support for heterogeneous graphs and node features, (3) they are not designed to generate plausible, realistic graphs as explanations. Although node feature support has been identified as important future work (Chen et al. 2023), most explainers still lack it. This is a critical gap, as GNNs tend to perform poorly without node features (Duong et al. 2019).
@@
In this paper, we introduce DiGNNExplainer, a model-level explanation approach for heterogeneous GNNs that synthesizes realistic heterogeneous graphs with node features (Figure 1). Our approach consists of five key steps: (1) We generate realistic graph structures using discrete denoising diffusion with DiGress (Vignac et al. 2023). (2) We generate realistic node features using discrete denoising diffusion. Specifically, we extend TabDDPM (Kotelnikov et al. 2023), a diffusion model originally designed for tabular data, to synthesize realistic, discrete node features—yielding a novel feature generator dubbed DiTabDDPM. Both our graph generator and our feature generator are trained on the same dataset as the GNN, keeping the generated graphs in-distribution, i.e., they reflect the patterns and statistical properties of the training data. (3) We check whether the synthesized graphs adhere to the metagraph and discard graphs that do not. (4) For each class, we select the synthetic graph with node features that maximizes the GNN’s prediction score for that class, ensuring that the selected graph captures the essential information relevant to the model’s decision-making. (5) We evaluate our synthesized explanation graphs with node features both in terms of realism and faithfulness. Specifically, we assess whether they preserve graph distributions, e.g., node degrees, clustering coefficients, and spectrum in terms of maximum mean discrepancy (MMD), and the generated explanations are faithful, i.e, achieve a high GNN prediction score (predictive faithfulness), and retain graph communities of the original graph (ground-truth faithfulness). We compare DiGNNExplainer to previous model-level explanation approaches and conduct an ablation study on different node feature generation techniques.
@@
  • Unlike existing tabular data generators (Kotelnikov et al. 2023; Jo, Lee, and Hwang 2022), which map all node features into a continuous space, our novel DiTabDDPM directly models discrete features using discrete diffusion, improving performance.
@@
explain why a GNN predicts a certain class by generating a graph that maximizes the predicted score for this class. Notable methods include XGNN (Yuan et al. 2020), GNNInterpreter (Wang and Shen 2023), and D4Explainer (Chen et al. 2023). We compare them in terms of their graph generation, node feature generation, and evaluation. Graph generation: XGNN employs reinforcement learning to iteratively add edges or nodes with one-hot encoded types. GNNInterpreter uses a probabilistic generative model generating edges and nodes independently with a fixed probability. D4Explainer, like our approach, utilizes diffusion-based graph generation but lacks support for different node types. Node feature generation: XGNN only supports one-hot encoded node types. GNNInterpreter proposes the use of actual node features but evaluates only node types. D4Explainer briefly discusses node features but focuses on structural graph generation. Evaluation: All three methods evaluate graphs based on predicted class scores. Only D4Explainer considers simple graph statistics; XGNN and GNNInterpreter do not. None of the approaches quantitatively assesses the graphs in terms of generated communities and distribution of generated node features. While some methods provide limited support for node features, none of them effectively address node feature generation as shown in our evaluation.
@@
Alternative explanation strategies use description logics or natural language instead of graph structures as explanation (Köhler and Heindorf 2024; Sapkota, Köhler, and Heindorf 2024; Heindorf and Neib 2025; Himmelhuber et al. 2021; Pluska et al. 2024), e.g., via OntoLearn (Demir et al. 2025) and EvoLearner (Heindorf et al. 2022). While some of these approaches support node features, they do not generate graphs as final explanations.
Instance-level GNN explainers
@@
such as GNNExplainer (Ying et al. 2019) and SubgraphX (Yuan et al. 2021) explain each GNN prediction by identifying a different subgraph, making it challenging to understand the overall GNN behavior. As pointed out by Chen et al. (2023), these methods derive their subgraphs from the training datasets, leading to unreliable explanations for out-of-distribution data.
Graph generators
@@
learn the distribution of a set of graphs to generate similar graphs. GraphVAE (Simonovsky and Komodakis 2018) employs a variational autoencoder to embed graphs into a continuous latent space and generates new graphs by sampling from this space. DiGress (Vignac et al. 2023) uses a discrete diffusion-based approach reconstructing graphs step-by-step in a discrete space, preserving structural information and node-type distributions more effectively. Both GraphVAE and DiGress have demonstrated strong performance on molecular datasets like MUTAG and QM9. HGEN (Ling, Yang, and Zhao 2021) employs a Generative Adversarial Network (GAN) to preserve metapath relations of heterogeneous datasets, such as DBLP, PubMed and IMDB. HGEN generates graphs with the same number of nodes as the input graph with a minimum number of 100 nodes, limiting its ability to produce small, human-intelligible graphs that may serve as an explanation. These approaches focus on structural graph generation with limited support for actual node features: HGEN does not generate node features at all, while GraphVAE and DiGress only generate discrete, categorical node features focusing on one-hot encoding of node types. Their performance for node feature generation has not been evaluated. Moreover, none of these approaches has been used and evaluated for generating explanation graphs as done in this work.
Tabular feature generation
@@
TabDDPM (Kotelnikov et al. 2023) uses diffusion models to synthesize discrete and continuous tabular data by learning a continuous space. To the best of our knowledge, it has never been used and evaluated to generate node features of graphs. Moreover, it has never been used for generating explanations. We extend TabDDPM to synthesize realistic discrete features by modeling them directly within a discrete space.
DiGNNExplainer
@@
Heterogeneous Graph with Node Features
@@
Generating Graphs and Node Features with Diffusion Models
We generate explanation graphs and node features through diffusion approaches (see Figure  1). We do so (1) using diffusion models for the graph structure and (2) diffusion models for the node features, (3) discard graphs that are inconsistent to the metagraph, and (4) we select one explanation graph per class. In the following, we give an overview of our approach. Detailed descriptions of the diffusion processes using DiGress and TabDDPM, and formal definitions of DiGNNExplainer, are provided in the Appendix.
@@ 1. Graph generation
We build upon DiGress (Vignac et al. 2023) to synthesize heterogeneous graph structures. Since DiGress is designed to generate graphs of a fixed size $n$ and requires training data of the same size, while the best size of an explanation graph is not known in advance, we create multiple training sets for DiGress with different graph sizes and train a different DiGress model for each graph size. [^1] In our evaluation, we typically use graph sizes between 10 and 15, as they provide meaningful explanations while remaining interpretable. For graph classification, we train the graph generators directly on datasets containing multiple small graphs of different sizes. In contrast, for node classification, where the data consists of a single large graph, we extract subgraphs of varying sizes using graph sampling techniques: We employ the forest fire sampling algorithm (Rozemberczki, Kiss, and Sarkar 2020), known for preserving graph connectivity and performing well in prior studies. To ensure that the sampled subgraphs include the nodes to be classified, we initiate the sampling process at nodes of the relevant types. For example, we start a fire at author nodes in the DBLP dataset. Specifically, for each graph size $n$, we generate $k$ subgraphs by repeating the following process until $k$ valid subgraphs are obtained: First, we randomly select a node of the type that is to be classified, then we start the fire from this node and stop once exactly $n$ nodes have been burned. If less than $n$ nodes are burned, we repeat the process. The nodes burned serve as sampled graphs. Following prior work, we use the default hyperparameters of the little-ball-of-fur library (Rozemberczki, Kiss, and Sarkar 2020), including a burning probability of 0.4.222https://little-ball-of-fur.readthedocs.io/en/latest/˙modules/littleballoffur/exploration˙sampling/forestfiresampler.html
@@ 2. Node feature generation
Given a node type $\tau$ with node feature matrix $\mathbf{X}_{\tau}$ (cf., Section Heterogeneous Graph with Node Features), we train multiple node feature generators as follows: If $\tau$ is the node type to be classified, we train one feature generator per class to simulate different feature distributions; if $\tau$ is not the node type to be classified, we train one feature generator; if $\tau$ is a node type with no node features at all, no feature generator is trained. For example, for the DBLP dataset, we train six node feature generators: four node feature generators for the four different classes of Author nodes (database, data mining, artificial intelligence, information retrieval), one for Paper nodes, one for Term nodes, and zero for conference nodes as they do not have node features.
If $\tau$ is a node type with a continuous feature matrix, we use the original TabDDPM (Kotelnikov et al. 2023) approach with continuous noise; for each node type $\tau$ with discrete feature matrix $\mathbf{X}_{\tau}$, we use our novel DiTabDDPM which combines the discrete noise from DiGress (Vignac et al. 2023) and the MLPBlock architecture from TabDDPM (Kotelnikov et al. 2023) as follows:
Forward diffusion process. The forward diffusion process for node features follows DiGress (Vignac et al. 2023), using discrete noise, marginal transitions, and a cosine noise scheduler.
Reverse diffusion process. In the reverse process, for each time step $t$ and node type $\tau$, given a discrete feature vector $\mathbf{x_{t}}\in\mathcal{X}_{\tau}^{d_{\tau}}$, we encode it via one-hot encoding as a vector $\mathbf{x_{t}^{(ohe)}}\in\mathcal{X}_{\tau}^{d_{\tau}\times k_{\tau}}$ into the denoising network which outputs the feature distribution vector $\mathbf{p}_{t-1}\in[0,1]^{d_{\tau}\times k_{\tau}}$ for its previous step. Then we sample the discrete features $\mathbf{x}_{t-1}^{(ohe)}\in\mathcal{X}_{\tau}^{d_{\tau}\times k_{\tau}}$ of the previous time step t1t-1 from the multinomial distribution defined by $\mathbf{p}_{t-1}$ and repeat the process.
Node feature loss. The node feature training loss $L_{train}$ for a single node is the cross entropy
$L_{train}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{d^{\tau}}\sum_{i=1}^{d_{\tau}}\mathrm{CrossEntropy}(\mathbf{p}_{pred},\mathbf{x}^{(i)}_{true})$
where $\mathbf{p}_{pred}\in[0,1]^{k_{\tau}}$ is the vector of predicted probabilities, and $\mathbf{x}^{(i)}_{true}\in\{x^{1}_{\tau},\ldots,x^{k_{\tau}}_{\tau}\}$ is the true discrete value of feature $i$ in the dataset.
@@
Evaluation
In this section, we evaluate how well our DiGNNExplainer explains the predictions of GNNs. After briefly introducing our evaluation setup, we compare DiGNNExplainer to three state-of-the-art GNN explanation methods. Additionally, we compare our discrete diffusion-based approach with a variant that uses variational autoencoders, and we investigate the impact of varying the number of discrete features. Our code is publicly available. [^3] https://github.com/ds-jrg/DiGNNExplainer Additional experimental results, including dataset statistics, hyperparameters, and implementation details, can be found in the supplementary material.
Evaluation Setup

### v2

@@ Abstract.
Graph neural networks; Explainable artificial intelligence; Diffusion models
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792518isbn: 979-8-4007-2307-0/2026/04ccs: Computing methodologies Supervised learning by classificationccs: Computing methodologies Neural networks
1. Introduction
@@
Figure 1. Overview of DiGNNExplainer. The example, based on the DBLP dataset, illustrates the process of generating explanations for a node classification task. Nodes represent different entity types: authors (red, $\bullet$), papers (blue, $\bullet$), conferences (orange, $\bullet$), and terms (green, $\bullet$). Node features, such as discrete bag-of-words representations of author keywords (e.g., [010100]), are encoded in red. The task is to predict the class of each author, i.e., their research area: database, data mining, artificial intelligence, or information retrieval. (1) & (2) The graph generator takes as input small real-world graphs with discrete node features and produces synthetic graphs with discrete node features. (3) The validity of these generated graphs is checked against the metagraph. (4) A trained GNN, which is to be explained, is applied to each of the nine author nodes of valid, synthesized graphs, producing a prediction for each of the four classes. For each class, we identify the node and corresponding graph with the highest prediction probability, yielding one explanation graph per class. (5) To evaluate our approach, we compute various metrics, including the ground-truth faithfulness between our explanation graphs and the communities within the real dataset.
@@
Complex real-world data, such as citation networks, social networks, and molecular structures, often exhibit an underlying graph structure. Nodes in these graphs typically have different types and are associated with additional features, i.e., they are heterogeneous graphs. For instance, in a citation network, there might be nodes of type “Paper” and “Author” each of which has additional node features like keywords or affiliations. Whereas related work sometimes uses the term “node feature” only to refer to a one-hot encoding of the node type (Yuan et al., 2020; Wang and Shen, 2023), we refer to actual attributes of nodes (e.g., age of persons, keywords of papers). A key machine learning task on such graphs is node classification, e.g., for fake news detection (Zhang et al., 2024a), corporate risk assessment (Zhang et al., 2024b), or predicting the mutagenicity of molecules (Luo et al., 2020). Towards this end, a number of powerful approaches for heterogeneous graphs have been proposed, e.g., Heterogeneous Graph Neural Networks (HGNNs) (Wang et al., 2019).
@@
However, the predictions made by HGNNs are opaque, providing no insight into the reasoning behind their decisions. In high-stakes applications—such as healthcare, finance, or criminal justice—this lack of transparency is a significant limitation, as it prevents users from understanding the rationale behind the model’s decisions. Towards this end, several post-hoc explanation methods have been proposed, such as GNNExplainer (Ying et al., 2019), SubgraphX (Yuan et al., 2021), PGExplainer (Luo et al., 2020), XGNN (Yuan et al., 2020), GNNInterpreter (Wang and Shen, 2023), and D4Explainer (Chen et al., 2023). Despite their contributions, these methods face significant limitations: (1) the first three do not provide model-level explanations, (2) the second three lack support for heterogeneous graphs and node features, (3) they are not designed to generate plausible, realistic graphs as explanations. Although node feature support has been identified as important future work (Chen et al., 2023), most explainers still lack it. This is a critical gap, as GNNs tend to perform poorly without node features (Duong et al., 2019).
@@
In this paper, we introduce DiGNNExplainer, a model-level explanation approach for heterogeneous GNNs that synthesizes realistic heterogeneous graphs with node features (Figure 1). Our approach consists of five key steps: (1) We generate realistic graph structures using discrete denoising diffusion with DiGress (Vignac et al., 2023). (2) We generate realistic node features using discrete denoising diffusion. Specifically, we extend TabDDPM (Kotelnikov et al., 2023), a diffusion model originally designed for tabular data, to synthesize realistic, discrete node features—yielding a novel feature generator dubbed DiTabDDPM. Both our graph generator and our feature generator are trained on the same dataset as the GNN, keeping the generated graphs in-distribution, i.e., they reflect the patterns and statistical properties of the training data. (3) We check whether the synthesized graphs adhere to the metagraph and discard graphs that do not. (4) For each class, we select the synthetic graph with node features that maximizes the GNN’s prediction score for that class, ensuring that the selected graph captures the essential information relevant to the model’s decision-making. (5) We evaluate our synthesized explanation graphs with node features both in terms of realism and faithfulness. Specifically, we assess whether they preserve graph distributions, e.g., node degrees, clustering coefficients, and spectrum in terms of maximum mean discrepancy (MMD), and the generated explanations are faithful, i.e., achieve a high GNN prediction score (predictive faithfulness), and retain graph communities of the original graph (ground-truth faithfulness). We compare DiGNNExplainer to previous model-level explanation approaches and conduct an ablation study on different node feature generation techniques.
@@
  • Unlike existing tabular data generators (Kotelnikov et al., 2023; Jo et al., 2022), which map all node features into a continuous space, our novel DiTabDDPM directly models discrete features using discrete diffusion, improving performance.
@@ 2. Related Work
Model-level GNN explainers explain why a GNN predicts a certain class by generating a graph that maximizes the predicted score for this class. Notable methods include XGNN (Yuan et al., 2020), GNNInterpreter (Wang and Shen, 2023), and D4Explainer (Chen et al., 2023). We compare them in terms of their graph generation, node feature generation, and evaluation. Graph generation: XGNN employs reinforcement learning to iteratively add edges or nodes with one-hot encoded types. GNNInterpreter uses a probabilistic generative model generating edges and nodes independently with a fixed probability. D4Explainer, like our approach, utilizes diffusion-based graph generation but lacks support for different node types. Node feature generation: XGNN only supports one-hot encoded node types. GNNInterpreter proposes the use of actual node features but evaluates only node types. D4Explainer briefly discusses node features but focuses on structural graph generation. Evaluation: All three methods evaluate graphs based on predicted class scores. Only D4Explainer considers simple graph statistics; XGNN and GNNInterpreter do not. None of the approaches quantitatively assesses the graphs in terms of generated communities and distribution of generated node features. While some methods provide limited support for node features, none of them effectively address node feature generation as shown in our evaluation.
@@
Alternative explanation strategies use description logics or natural language instead of graph structures as explanation (Köhler and Heindorf, 2024; Sapkota et al., 2024; Heindorf and Neib, 2025; Himmelhuber et al., 2021; Pluska et al., 2024), e.g., via OntoLearn (Demir et al., 2025) and EvoLearner (Heindorf et al., 2022). While some of these approaches support node features, they do not generate graphs as final explanations.
@@
Instance-level GNN explainers such as GNNExplainer (Ying et al., 2019) and SubgraphX (Yuan et al., 2021) explain each GNN prediction by identifying a different subgraph, making it challenging to understand the overall GNN behavior. As pointed out by Chen et al. (2023), these methods derive their subgraphs from the training datasets, leading to unreliable explanations for out-of-distribution data.
@@
Graph generators learn the distribution of a set of graphs to generate similar graphs. GraphVAE (Simonovsky and Komodakis, 2018) employs a variational autoencoder to embed graphs into a continuous latent space and generates new graphs by sampling from this space. DiGress (Vignac et al., 2023) uses a discrete diffusion-based approach reconstructing graphs step-by-step in a discrete space, preserving structural information and node-type distributions more effectively. Both GraphVAE and DiGress have demonstrated strong performance on molecular datasets like MUTAG and QM9. HGEN (Ling et al., 2021) employs a Generative Adversarial Network (GAN) to preserve metapath relations of heterogeneous datasets, such as DBLP, PubMed, and IMDB. HGEN generates graphs with the same number of nodes as the input graph with a minimum number of 100 nodes, limiting its ability to produce small, human-intelligible graphs that may serve as an explanation. These approaches focus on structural graph generation with limited support for actual node features: HGEN does not generate node features at all, while GraphVAE and DiGress only generate discrete, categorical node features focusing on one-hot encoding of node types. Their performance for node feature generation has not been evaluated. Moreover, none of these approaches has been used and evaluated for generating explanation graphs as done in this work.
@@
Tabular feature generation TabDDPM (Kotelnikov et al., 2023) uses diffusion models to synthesize discrete and continuous tabular data by learning a continuous space. To the best of our knowledge, it has never been used and evaluated to generate node features of graphs. Moreover, it has never been used for generating explanations. We extend TabDDPM to synthesize realistic discrete features by modeling them directly within a discrete space.
3. DiGNNExplainer
@@
3.1. Heterogeneous Graph with Node Features
@@
3.2. Generating Graphs and Node Features with Diffusion Models
We generate explanation graphs and node features through diffusion approaches (see Figure  1). We do so (1) using diffusion models for the graph structure and (2) diffusion models for the node features, (3) discard graphs that are inconsistent with the metagraph, and (4) we select one explanation graph per class. In the following, we give an overview of our approach. Detailed descriptions of the diffusion processes using DiGress and TabDDPM, and formal definitions of DiGNNExplainer, are provided in the Appendix.
@@ 1. Graph generation
We build upon DiGress (Vignac et al., 2023) to synthesize heterogeneous graph structures. Since DiGress performs well in generating graphs of a fixed size $n$ and requires training data of the same size, while the best size of an explanation graph is not known in advance, we create multiple training sets for DiGress with different graph sizes and train a different DiGress model for each graph size. [^1] In our evaluation, we typically use graph sizes between 10 and 15, as they provide meaningful explanations while remaining interpretable. For graph classification, we train the graph generators directly on datasets containing multiple small graphs of different sizes. In contrast, for node classification, where the data consists of a single large graph, we extract subgraphs of varying sizes using graph sampling techniques: We employ the forest fire sampling algorithm (Rozemberczki et al., 2020), known for preserving graph connectivity and performing well in prior studies. To ensure that the sampled subgraphs include the nodes to be classified, we initiate the sampling process at nodes of the relevant types. For example, we start a fire at author nodes in the DBLP dataset. Specifically, for each graph size $n$, we generate $k$ subgraphs by repeating the following process until $k$ valid subgraphs are obtained: First, we randomly select a node of the type that is to be classified, then we start the fire from this node and stop once exactly $n$ nodes have been burned. If fewer than $n$ nodes are burned, we repeat the process. The nodes burned serve as sampled graphs. Following prior work, we use the default hyperparameters of the little-ball-of-fur library (Rozemberczki et al., 2020), including a burning probability of 0.4.222https://little-ball-of-fur.readthedocs.io/en/latest/_modules/littleballoffur/exploration_sampling/forestfiresampler.html
@@ 2. Node feature generation
Given a node type $\tau$ with node feature matrix $\mathbf{X}_{\tau}$ (cf., Section 3.1), we train multiple node feature generators as follows: If $\tau$ is the node type to be classified, we train one feature generator per class to simulate different feature distributions; if $\tau$ is not the node type to be classified, we train one feature generator; if $\tau$ is a node type with no node features at all, no feature generator is trained. For example, for the DBLP dataset, we train six node feature generators: four node feature generators for the four different classes of author nodes (database, data mining, artificial intelligence, information retrieval), one for paper nodes, one for term nodes, and zero for conference nodes as they do not have node features.
If $\tau$ is a node type with a continuous feature matrix, we use the original TabDDPM (Kotelnikov et al., 2023) approach with continuous noise; for each node type $\tau$ with discrete feature matrix $\mathbf{X}_{\tau}$, we use our novel DiTabDDPM which combines the discrete noise from DiGress (Vignac et al., 2023) and the MLPBlock architecture from TabDDPM (Kotelnikov et al., 2023) as follows:
Forward diffusion process. The forward diffusion process for node features follows DiGress (Vignac et al., 2023), using discrete noise, marginal transitions, and a cosine noise scheduler.
Reverse diffusion process. In the reverse process, for each time step $t$ and node type $\tau$, given a discrete feature vector $\mathbf{x_{t}}\in\mathcal{X}_{\tau}^{d_{\tau}}$, we encode it via one-hot encoding as a vector $\mathbf{x_{t}^{(ohe)}}\in\{0,1\}^{d_{\tau}\times k_{\tau}}$ into the denoising network which outputs the feature distribution vector $\mathbf{p}_{t-1}\in[0,1]^{d_{\tau}\times k_{\tau}}$ for its previous step. Then we sample the discrete features $\mathbf{x}_{t-1}\in\mathcal{X}_{\tau}^{d_{\tau}}$ of the previous time step t1t-1 from the multinomial distribution defined by $\mathbf{p}_{t-1}$ and repeat the process.
Node feature loss. The node feature training loss $L_{train}$ for a single node is the cross-entropy
$L_{train}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{d_{\tau}}\sum_{i=1}^{d_{\tau}}\mathrm{CrossEntropy}(\mathbf{p}^{(i)}_{pred},\mathbf{x}^{(i)}_{true})$
where $\mathbf{p}^{(i)}_{pred}\in[0,1]^{k_{\tau}}$ is the vector of predicted probabilities, and $\mathbf{x}^{(i)}_{true}\in\{x^{1}_{\tau},\ldots,x^{k_{\tau}}_{\tau}\}$ is the true discrete value of feature $i$ in the dataset.
@@
4. Evaluation
In this section, we evaluate how well our DiGNNExplainer explains the predictions of GNNs. After briefly introducing our evaluation setup, we compare DiGNNExplainer to three state-of-the-art GNN explanation methods. Additionally, we compare our discrete diffusion-based approach with a variant that uses variational autoencoders, and we investigate the impact of varying the number of discrete features. Our code is publicly available. [^3] https://github.com/ds-jrg/DiGNNExplainer Additional experimental results, including dataset statistics, hyperparameters, and implementation details, can be found in the Appendix.
4.1. Evaluation Setup